# Towards Vision-Language Geo-Foundation Model: A Survey

### v1

@@ 3.2.3 Generative VLGFMs
A typical generative VLGFM comprises three main components: a text encoder, a pre-trained conditional diffusion model, and an image decoder. A schematic of this architecture is depicted in Figure 2. The working mechanism of diffusion models consists of two phases. Initially, noise is introduced into the dataset, which is a crucial step in the forward diffusion process, followed by a systematic reversal of this process. In other words, the method involves identifying a stochastic differential equation (SDE) [83] whose process converges to a standard normal distribution. This equation transforms the data, mapping it onto the standard normal distribution. Subsequently, samples that conform to the original data distribution are generated by sampling from this distribution and applying the inverse of the SDE to the sampled data.
Conditional diffusion models are a class of generative models that extend the fundamental principles of diffusion processes to incorporate conditional information, thereby enabling the generation of data that adheres to specific conditions or constraints. In the field of remote sensing, DiffusionSat[37] successfully enabled the model to generate satellite images of specific locations and seasons by incorporating information such as latitude, longitude, and the time of capture during the training process of the diffusion model. Furthermore, CRS-Diff [74] attempted to guide the diffusion model in generating RSIs with similar layouts by using binary masks or road sketches, which can be used to generate samples that are difficult to collect in the real world. Due to limitations in the number of training samples and computational resources, the work on generative models in the domain of remote sensing is still relatively scarce.
@@ 3.3 Capabilities
Inspired by MMBench [59], we have categorized the capabilities exhibited by VLGFMs into three hierarchical levels and clarified the design principles underlying this classification.
@@
Overview. Figure 3 provides an overview of the existing multimodal capabilities of VLGFM. This benchmark encompasses 20 different leaf capabilities, offering a comprehensive evaluation of various fine-grained perceptual and reasoning abilities. Figure 4 illustrates each capability through specific examples, while Table V provides the capabilities supported by major remote sensing vision-language benchmarks. Since certain capabilities currently lack corresponding qualitative analysis datasets, they are not displayed in this table. Please note that the classification displayed is not the final version. We plan to expand the capability taxonomy further to make it more comprehensive.
@@
The Hierarchical Taxonomy of Capability. Humans possess remarkable perceptual and reasoning abilities, enabling them to understand and interact with the world. These capabilities have been crucial in human evolution and are the foundation for complex cognitive processes. Perception involves gathering information from sensory inputs, while reasoning involves drawing conclusions based on this information. Collectively, they underpin the foundation of most tasks in the real world, including object recognition, problem-solving, and decision-making [84, 85]. VLGFMs are also expected to exhibit strong perceptual and reasoning capabilities in pursuit of artificial intelligence for remote sensing verticals. Therefore, our capability taxonomy designates perception and reasoning as our top-level capability dimensions, referred to as L-1 capability dimensions. For L-2 capabilities, we derive image-wise perception, region-wise perception, instance-wise perception, pixel-wise perception from L-1 perception, Attribute reasoning, and logical reasoning from L-1 reasoning. To make our benchmark as granular as possible, we further derive 20 L-3 capability dimensions from L-2.
@@ 3.3.1 Perception
Perceptual capability is the foundational skill for interpreting RSIs. With the rise of deep learning technology, numerous specialized models have been proposed to address image perception tasks in remote sensing, such as scene classification [38], object detection [13], and segmentation [86]. The emergence of VLGFM provides a solution for handling all perception tasks using a single model. Perception can be mainly divided into three types: image-wise, instance-wise, and pixel-wise perception.
@@
Imgae-wise perception belongs to coarse-grained perception and is relatively easier. It can be divided into image scene, image quality, image modality, image captioning, and image retrieval. Instance-wise perception belongs to fine-grained perception, which requires identifying objects within an image and is more challenging. It can be divided into eight categories, encompassing some visual grounding capabilities. The basic building blocks of an image are pixels; thus, pixel-wise perception represents the most fine-grained perceptual capability. Pixel-wise perception often requires an additional pixel decoder. The SAM decoder [73] is a common choice. The output features of VLGFM will be aligned in dimensions and then concatenated with the prompt features of SAM. Finally, they are fed into the SAM decoder to generate the mask. Pixel-wise perception can be divided into referring expression segmentation and open-vocabulary segmentation.
@@
$\bullet$ Image Scene aims to categorize an image into a predefined set of scenes or environments based on its overall visual content. This review, referencing the SATIN framework [38], categorizes remote sensing scenes into six distinct categories. Land cover refers to the biophysical surface characteristics of the Earth, with examples including forests, grasslands, and buildings. Land use describes the economic and social functions of areas, such as residential zones, industrial sectors, and parking lots, including bodies of water utilized for specific purposes. Hierarchical land use takes this a step further by classifying land use into more detailed levels of granularity, for instance, categorizing an area broadly as a transportation area, then more specifically as highway land, and even more precisely as a bridge. Complex scenes are images that contain one or more land use labels, indicating a blend of different uses or covers in one scene. Rare scenes represent unusual or infrequent categories that do not typically fall under standard land cover or land use definitions, such as hurricane damage or wind turbines. Lastly, false color scenes are images represented in false color, focusing on a single land use or land cover class, aiding in the distinct visualization of specific area characteristics.
@@
$\bullet$ Image Quality aims to evaluate the condition of RSIs, taking into account factors such as imaging quality, noise levels [87], cloud cover [88], and low-light conditions [89]. This capability helps users filter out low-quality RSIs to prevent noise from contaminating the results. To date, no remote sensing benchmarks are specifically designed to test image quality capabilities.
@@
$\bullet$ Image Captioning aims to produce a coherent and contextually relevant narrative in natural language that encapsulates the overall scene depicted in an image. Image captioning can further be categorized based on the input type into change captioning [90], multi-temporal image captioning, and video captioning [91]. The input for the change captioning task consists of two images. The objective is to summarize and describe the changes observed in the later image relative to the earlier one in the temporal sequence. Compared to change captioning, which involves two images, multi-temporal image captioning deals with a larger set of images and aims to encapsulate the progression of changes exhibited across a series of temporal images. In the case of video captioning, the input is a video sequence, and the task is to describe the events occurring within the video.
@@
$\bullet$ Change Detection aims to identify areas or objects that have changed in RSIs of the same place at different times. Compared to change captioning capabilities, it focuses more on specific instances in the image than the entire image. Observing changes in specific targets is a more practical function in real-world applications. Levir-CC [92] and CDVQA [93] have analyzed changes in vegetation and buildings, which are used to track the progression of urbanization.
$\bullet$ Object Counting aims to identify and count the instances of specific objects within an image. Due to the high density of targets typically present in RSIs, manual techniques are time-consuming and labour-intensive. This makes counting the number of targets in RSIs a particularly valuable capability. For example, we can count animals, trees, buildings, and vehicles in RSIs [94].
@@ 3.3.2 Reasoning
However, perception tasks can be accomplished using pure vision specifical models, while autonomous reasoning is the capability most eagerly anticipated in the field of remote sensing interpretation [95]. Reasoning poses greater difficulty and has a higher application value than perception. Reasoning can be mainly divided into two types: logic and attribute reasoning.
@@
$\bullet$ Geospatial Localization aims to pinpoint the geographic location represented in a remote sensing image, potentially identifying the specific continent, country, city, or precise latitude and longitude based on terrain and landmark features [96]. Its inspiration comes from the popular geography game GeoGuessr [^4] https://www.geoguessr.com, in which players are placed in a street view panorama somewhere in the world and must guess their location. Geospatial Localization typically necessitates acquiring information on topography, vegetation, and building styles through perceptual capabilities to perform reasoning, which adds significant complexity.
$\bullet$ Geometric Measurement aims to measure the physical dimensions of objects and features on the ground using RSIs, including estimating height, distance, and calculating area. The application of this capability is extensive; for example, it can be used to estimate the distance between two buildings in an image, the relative elevation from the foot to the summit of a mountain, the length of the ship, and the area occupied by cropland [97].
$\bullet$ Resource Assessment aims to estimate and analyze natural resource quantity, quality, and spatial distribution, such as estimating urban population sizes [95], calculating forest carbon dioxide emissions, and estimating crop yields.
@@
TABLE V: Capabilities supported by major remote sensing vision-language benchmarks. Abbreviations IS, IC, IR, CD, OP, OC, SR, AR, REC, REG, RES, OVS, GM and PP in the table stand for image scene, image captioning, image retrieval, change detection, object presence, object counting, spatial relationship, attribute recognition, open-vocabulary detection, referring expression comprehension, referring expression generation, referring expression segmentation, open-vocabulary segmentation, geospatial localization, geometric measurement, resource assessment, sensor property, and physical property, respectively.
Dataset IS IC IR CD OP OC SR AR REC REG RES OVS GM PP Total
Sydney-Captions [40] 1
UCM-Captions [40] 2
RSICD [43] 2
RSVQA-LR [41] 5
RSVQA-HR [41] 6
RSVQAxBEN [98] 2
FloodNet [99] 2
RSITMD [39] 2
RSIVQA [100] 4
NWPU-Captions [44] 1
LEVIR-CC [92] 2
CDVQA [93] 4
UAV [46] 1
RSVG [42] 5
CRSVQA [101] 5
DIOR-RSVG [47] 4
RSIEval [24] 6
SATIN [38] 1
NAIP-OSM [26] 2
RemoteCount [22] 1
SkyScript [27] 2
EarthVQA [102] 5
RRSIS [103] 1
GeoChat-Bench [25] 6
RRSIS-D [104] 1
VLEO-Bench [94] 6
FineGrip [105] 2
@@ 4.1 Challenges
$\bullet$ Training cost too high. As shown in Table I, VLGFMs have high hardware resource requirements, with most work completed using A100 GPUs. However, for some university laboratories, the cost of A100 GPUs is prohibitively high. LMDeploy [106] is attempting to reduce memory requirements during inference through quantization. However, the high demand for GPUs during the training phase inevitably raises the barrier for research in this direction.
$\bullet$ Benchmark too limited. On one hand, as shown in Table V, the existing benchmarks do not comprehensively cover various capabilities, making it difficult to measure the performance of VLGFMs fully. On the other hand, the tasks in existing multimodal benchmarks are relatively simple, with most humans being able to solve them easily. Hence, even if VLGFMs achieve high scores on these benchmarks, it is still not convincing.
$\bullet$ Correction of hallucinations. LLMs tend to generate absurd or untrue content unrelated to any sources, a phenomenon researchers call hallucination [107]. Due to the presence of hallucinations, current large models are still not widely applicable in high-precision tasks such as critical target interpretation and remote sensing mapping. To mitigate the hallucination phenomenon, researchers train models to refuse to answer questions they cannot handle. However, some researchers have a different perspective on the hallucination phenomenon. They argue that hallucination is not only a problem but also a potential capability. This capability can help LLMs generate unexpectedly interesting answers. Whether considered a capability or a challenge, the hallucination phenomenon in LLMs is an issue that must be addressed as we work to enhance model capabilities and practical applications.
@@ 4.2 Future Work
$\bullet$ Training-free technology. Given that the high training costs raise the barrier for VLGFMs, achieving good results in the remote sensing without extensive training would be a welcome development for researchers. This approach is referred to as training-free technology. Chain of thought (CoT) [108] might be one of the more promising training-free approaches. In this method, the model generates a reasoning process before providing an answer, similar to how humans often make correct judgments after careful consideration. Zero-shot CoT [109] employs a single, task-agnostic ”general” trigger to prompt the model to generate a reasoning chain. In the first stage, a phrase such as ”Let’s think step by step” is appended to the question and input into the LLM, prompting it to generate a step-by-step logical reasoning process for the problem. Next, the prompt from the first stage and the reasoning chain generated by the model are concatenated and used as the input for the second stage, where the model is then asked to generate the answer. However, this capability to generate logical chains only emerges when the model reaches a certain scale. It is worth mentioning that the aforementioned methods do not require training, which significantly lowers the research barrier.

### v2

@@ 3.2.3 Generative VLGFMs
A typical generative VLGFM comprises three main components: a text encoder, a pre-trained conditional diffusion model, and an image decoder. A schematic of this architecture is depicted in Figure 2. The working mechanism of diffusion models consists of two phases. Initially, noise is introduced into the dataset, which is a crucial step in the forward diffusion process, followed by a systematic reversal of this process. In other words, the method involves identifying a stochastic differential equation (SDE) song2021scorebased whose process converges to a standard normal distribution. This equation transforms the data, mapping it onto the standard normal distribution. Subsequently, samples that conform to the original data distribution are generated by sampling from this distribution and applying the inverse of the SDE to the sampled data.
Conditional diffusion models are a class of generative models that extend the fundamental principles of diffusion processes to incorporate conditional information, thereby enabling the generation of data that adheres to specific conditions or constraints. In the field of remote sensing, DiffusionSatkhanna2023diffusionsat successfully enabled the model to generate satellite images of specific locations and seasons by incorporating information such as latitude, longitude, and the time of capture during the training process of the diffusion model. Furthermore, CRS-Diff tang2024crsdiff attempted to guide the diffusion model in generating RSIs with similar layouts by using binary masks or road sketches, which can be used to generate samples that are difficult to collect in the real world. Due to limitations in the number of training samples and computational resources, the work on generative models in the domain of remote sensing is still relatively scarce.
@@ 3.3 Capabilities
Inspired by MMBench liu2023mmbench, we have categorized the capabilities exhibited by VLGFMs into three hierarchical levels and clarified the design principles underlying this classification.
@@
Overview. Figure 3 provides an overview of the existing multimodal capabilities of VLGFM. This benchmark encompasses 20 different leaf capabilities, offering a comprehensive evaluation of various fine-grained perceptual and reasoning abilities. Figure 4 illustrates each capability through specific examples, while Table 5 provides the capabilities supported by major remote sensing vision-language benchmarks. Since certain capabilities currently lack corresponding qualitative analysis datasets, they are not displayed in this table. Please note that the classification displayed is not the final version. We plan to expand the capability taxonomy further to make it more comprehensive.
@@
The Hierarchical Taxonomy of Capability. Humans possess remarkable perceptual and reasoning abilities, enabling them to understand and interact with the world. These capabilities have been crucial in human evolution and are the foundation for complex cognitive processes. Perception involves gathering information from sensory inputs, while reasoning involves drawing conclusions based on this information. Collectively, they underpin the foundation of most tasks in the real world, including object recognition, problem-solving, and decision-making fodor1983modularity; oaksford2007bayesian. VLGFMs are also expected to exhibit strong perceptual and reasoning capabilities in pursuit of artificial intelligence for remote sensing verticals. Therefore, our capability taxonomy designates perception and reasoning as our top-level capability dimensions, referred to as L-1 capability dimensions. For L-2 capabilities, we derive image-wise perception, region-wise perception, instance-wise perception, pixel-wise perception from L-1 perception, Attribute reasoning, and logical reasoning from L-1 reasoning. To make our benchmark as granular as possible, we further derive 20 L-3 capability dimensions from L-2.
@@ 3.3.1 Perception
Perceptual capability is the foundational skill for interpreting RSIs. With the rise of deep learning technology, numerous specialized models have been proposed to address image perception tasks in remote sensing, such as scene classification roberts2023satin, object detection mmrotate, and segmentation rsi-seg. The emergence of VLGFM provides a solution for handling all perception tasks using a single model. Perception can be mainly divided into three types: image-wise, instance-wise, and pixel-wise perception.
@@
Imgae-wise perception belongs to coarse-grained perception and is relatively easier. It can be divided into image scene, image quality, image modality, image captioning, and image retrieval. Instance-wise perception belongs to fine-grained perception, which requires identifying objects within an image and is more challenging. It can be divided into eight categories, encompassing some visual grounding capabilities. The basic building blocks of an image are pixels; thus, pixel-wise perception represents the most fine-grained perceptual capability. Pixel-wise perception often requires an additional pixel decoder. The SAM decoder SAM is a common choice. The output features of VLGFM will be aligned in dimensions and then concatenated with the prompt features of SAM. Finally, they are fed into the SAM decoder to generate the mask. Pixel-wise perception can be divided into referring expression segmentation and open-vocabulary segmentation.
@@
$\bullet$ Image Scene aims to categorize an image into a predefined set of scenes or environments based on its overall visual content. This review, referencing the SATIN framework roberts2023satin, categorizes remote sensing scenes into six distinct categories. Land cover refers to the biophysical surface characteristics of the Earth, with examples including forests, grasslands, and buildings. Land use describes the economic and social functions of areas, such as residential zones, industrial sectors, and parking lots, including bodies of water utilized for specific purposes. Hierarchical land use takes this a step further by classifying land use into more detailed levels of granularity, for instance, categorizing an area broadly as a transportation area, then more specifically as highway land, and even more precisely as a bridge. Complex scenes are images that contain one or more land use labels, indicating a blend of different uses or covers in one scene. Rare scenes represent unusual or infrequent categories that do not typically fall under standard land cover or land use definitions, such as hurricane damage or wind turbines. Lastly, false color scenes are images represented in false color, focusing on a single land use or land cover class, aiding in the distinct visualization of specific area characteristics.
@@
$\bullet$ Image Quality aims to evaluate the condition of RSIs, taking into account factors such as imaging quality, noise levels alamri2010noise, cloud cover SHEN2014cloud, and low-light conditions lowlight. This capability helps users filter out low-quality RSIs to prevent noise from contaminating the results. To date, no remote sensing benchmarks are specifically designed to test image quality capabilities.
@@
$\bullet$ Image Captioning aims to produce a coherent and contextually relevant narrative in natural language that encapsulates the overall scene depicted in an image. Image captioning can further be categorized based on the input type into change captioning 10271701, multi-temporal image captioning, and video captioning rs15082139. The input for the change captioning task consists of two images. The objective is to summarize and describe the changes observed in the later image relative to the earlier one in the temporal sequence. Compared to change captioning, which involves two images, multi-temporal image captioning deals with a larger set of images and aims to encapsulate the progression of changes exhibited across a series of temporal images. In the case of video captioning, the input is a video sequence, and the task is to describe the events occurring within the video.
@@
$\bullet$ Change Detection aims to identify areas or objects that have changed in RSIs of the same place at different times. Compared to change captioning capabilities, it focuses more on specific instances in the image than the entire image. Observing changes in specific targets is a more practical function in real-world applications. Levir-CC levir-cc and CDVQA cdvqa have analyzed changes in vegetation and buildings, which are used to track the progression of urbanization.
$\bullet$ Object Counting aims to identify and count the instances of specific objects within an image. Due to the high density of targets typically present in RSIs, manual techniques are time-consuming and labour-intensive. This makes counting the number of targets in RSIs a particularly valuable capability. For example, we can count animals, trees, buildings, and vehicles in RSIs vleobench.
@@ 3.3.2 Reasoning
However, perception tasks can be accomplished using pure vision specifical models, while autonomous reasoning is the capability most eagerly anticipated in the field of remote sensing interpretation roberts2024charting. Reasoning poses greater difficulty and has a higher application value than perception. Reasoning can be mainly divided into two types: logic and attribute reasoning.
@@
$\bullet$ Geospatial Localization aims to pinpoint the geographic location represented in a remote sensing image, potentially identifying the specific continent, country, city, or precise latitude and longitude based on terrain and landmark features NEURIPS2023_1b57aadd. Its inspiration comes from the popular geography game GeoGuessr [^4] https://www.geoguessr.com, in which players are placed in a street view panorama somewhere in the world and must guess their location. Geospatial Localization typically necessitates acquiring information on topography, vegetation, and building styles through perceptual capabilities to perform reasoning, which adds significant complexity.
$\bullet$ Geometric Measurement aims to measure the physical dimensions of objects and features on the ground using RSIs, including estimating height, distance, and calculating area. The application of this capability is extensive; for example, it can be used to estimate the distance between two buildings in an image, the relative elevation from the foot to the summit of a mountain, the length of the ship, and the area occupied by cropland roberts2023gpt4geo.
$\bullet$ Resource Assessment aims to estimate and analyze natural resource quantity, quality, and spatial distribution, such as estimating urban population sizes roberts2024charting, calculating forest carbon dioxide emissions, and estimating crop yields.
@@
Table 5: Capabilities supported by major remote sensing vision-language benchmarks. Abbreviations IS, IC, IR, CD, OP, OC, SR, AR, REC, REG, RES, OVS, GM and $P$ in the table stand for image scene, image captioning, image retrieval, change detection, object presence, object counting, spatial relationship, attribute recognition, open-vocabulary detection, referring expression comprehension, referring expression generation, referring expression segmentation, open-vocabulary segmentation, geospatial localization, geometric measurement, resource assessment, sensor property, and physical property, respectively.
Dataset IS IC IR CD OP OC SR AR REC REG RES OVS GM $P$ Total
Sydney-Captions sydneycaption 1
UCM-Captions sydneycaption 2
RSICD rsicd 2
RSVQA-LR rsvqa 5
RSVQA-HR rsvqa 6
RSVQAxBEN rsvqaben 2
FloodNet floodnet 2
RSITMD rsitmd 2
RSIVQA rsivqa 4
NWPU-Captions nwpucaption 1
LEVIR-CC levir-cc 2
CDVQA cdvqa 4
UAV uavcaption 1
RSVG rsvg 5
CRSVQA crsvqa 5
DIOR-RSVG dior-rsvg 4
RSIEval hu2023rsgpt 6
SATIN roberts2023satin 1
NAIP-OSM mall2023remote 2
RemoteCount liu2024remoteclip 1
SkyScript wang2024skyscript 2
EarthVQA wang2024earthvqa 5
RRSIS rrsis 1
GeoChat-Bench kuckreja2023geochat 6
RRSIS-D rrsis-d 1
ChatEarthNet yuan2024chatearthnet 3
VLEO-Bench vleobench 6
FineGrip finegrip 2
RS-GPT4V xu2024rsgpt4v 6
VRSBench li2024vrsbench 7
@@ 4.1 Challenges
$\bullet$ Training cost too high. As shown in Table 1, VLGFMs have high hardware resource requirements, with most work completed using A100 GPUs. However, for some university laboratories, the cost of A100 GPUs is prohibitively high. LMDeploy 2023lmdeploy is attempting to reduce memory requirements during inference through quantization. However, the high demand for GPUs during the training phase inevitably raises the barrier for research in this direction.
$\bullet$ Benchmark too limited. On one hand, as shown in Table 5, the existing benchmarks do not comprehensively cover various capabilities, making it difficult to measure the performance of VLGFMs fully. On the other hand, the tasks in existing multimodal benchmarks are relatively simple, with most humans being able to solve them easily. Hence, even if VLGFMs achieve high scores on these benchmarks, it is still not convincing. Furthermore, zhu2024foundation defines eleven key features of geo-foundational models that are beneficial for geoscientific downstream applications, which complement this survey.
$\bullet$ Correction of hallucinations. LLMs tend to generate absurd or untrue content unrelated to any sources, a phenomenon researchers call hallucination guan2024hallusionbench. Due to the presence of hallucinations, current large models are still not widely applicable in high-precision tasks such as critical target interpretation and remote sensing mapping. To mitigate the hallucination phenomenon, researchers train models to refuse to answer questions they cannot handle. However, some researchers have a different perspective on the hallucination phenomenon. They argue that hallucination is not only a problem but also a potential capability. This capability can help LLMs generate unexpectedly interesting answers. Whether considered a capability or a challenge, the hallucination phenomenon in LLMs is an issue that must be addressed as we work to enhance model capabilities and practical applications.
@@ 4.2 Future Work
$\bullet$ Training-free technology. Given that the high training costs raise the barrier for VLGFMs, achieving good results in the remote sensing without extensive training would be a welcome development for researchers. This approach is referred to as training-free technology. Chain of thought (CoT) NEURIPS2022_CoT might be one of the more promising training-free approaches. In this method, the model generates a reasoning process before providing an answer, similar to how humans often make correct judgments after careful consideration. Zero-shot CoT NEURIPS2022_zeroshotCoT employs a single, task-agnostic ”general” trigger to prompt the model to generate a reasoning chain. In the first stage, a phrase such as ”Let’s think step by step” is appended to the question and input into the LLM, prompting it to generate a step-by-step logical reasoning process for the problem. Next, the prompt from the first stage and the reasoning chain generated by the model are concatenated and used as the input for the second stage, where the model is then asked to generate the answer. However, this capability to generate logical chains only emerges when the model reaches a certain scale. It is worth mentioning that the aforementioned methods do not require training, which significantly lowers the research barrier.
@@ 5 Conclusions
6 Abbreviations
CoT: chain of thought; Det.: detection; GEE: Google Earth Engine; GM: geometric measurement; GSD: ground sampling distance; IC: image captioning; IR: image retrieval; IS: image scene classification; LLMs: large language models; OC: object counting; OSM: OpenStreetMap; OVD: open-vocabulary detection; REC: referring expression comprehension; REG: referring expression generation; RES: referring expression segmentation; RSIs: remote sensing images; Seg.: segmentation; SDE: stochastic differential equation; VC: video captioning; VGFMs: vision geo-foundation models; VG: visual grounding; VLFMs: vision-language foundation models; VLGFMs: vision-language geo-foundation models; VQA: visual question answering.
Data Availability: The datasets generated during and/or analyzed during the current study are available in the Awesome-VLGFM repository, https://github.com/zytx121/Awesome-VLGFM.
Competing Interests: The authors have no relevant financial or non-financial interests to disclose.
Authors’ Contributions: Idea: Yue Zhou; Conceptualization: Xue Yang; Literature Search: Yue Zhou, Xue Yang; Data Analysis Yue Zhou; Writing - original draft preparation: Yue Zhou; Writing - review and editing: Xue Yang; Funding acquisition: Xue Yang; Supervision: Xue Yang.
Funding: This work was supported by National Natural Science Foundation of China (62506229), and Natural Science Foundation of Shanghai (25ZR1402268).
Acknowledgements.
Acknowledgements: Not applicable.